# Invariant Subspaces of Perturbed Backward Shift

### v1

@@ 3. Perturbations of $T_{z}^{*}$
The setting of Theorem 2.1 is exceedingly broad, embracing invariant subspaces of finite rank perturbations of inner function-based Toeplitz operators. This one-sided result fits Hitt and Sarason’s perspectives. Indeed, their results were only one-directional. In this section, we depart from this pattern and try to explain a converse direction. Our converse, however, makes better sense for the particular case of the special inner function: $\varphi(z)=z$, $z\in\mathbb{D}$. We keep on assuming that $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets of vectors in $H^{2}$.
To address the reverse direction, we first note a consequence of Corollary 2.2: There exists a unitary operator $\mathcal{J}:\mathcal{K}\to\mathcal{M}$, such that
@@
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$. This readily follows from the fact that an element $(F,f_{1})\in\mathcal{K}_{\Psi}$ can be written as $(F,f_{0})=(k_{1},\ldots,k_{p},k_{p+1})$ (here we are rewriting $f_{0}=k_{p+1}$ for simplicity), so that
@@
We also have the identity that $\|F\|^{2}+\|f_{0}\|^{2}=\|(F,f_{0})\|^{2}$.
We fix a rank-$m$ perturbation of $T_{z}^{*}$ as $T_{z}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, and set
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}u_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ such that
@@
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|{F}\|^{2}+\|{f_{0}}\|^{2}=\|F_{0}F+f_{0}\|^{2}\},$ (3.3)
@@
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$. Conversely, if $\mathcal{M}$ has the representation (3.2) for some orthonormal basis $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ with $\mathcal{K}_{\Psi}$ as in (3.3), and a unitary $\mathcal{J}$ as in (3.4), then $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i}$.
@@
=$\mathcal{M}=\{f\in H^{2}:f=F_{0}F+f_{0}\text{ with }\|f\|^{2}=\|F\|^{2}+\|{f_{0}}\|^{2}\text{ for some }(F,f_{0})\in\mathcal{K}_{\Psi}\}.$
@@
$(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i})f$ $=[T_{z}^{*}-\sum_{i=1}^{p}\langle\cdot,f_{i}\rangle T_{z}^{*}f_{i}](F_{0}F+f_{0})$
@@
for all $i=1,\ldots,p$, where $e_{i}=(0,\ldots,0,1,0,\ldots,0)$ with $1$ at the $i$-th position and zero elsewhere. This conclusion arises immediately from the identity (3.1). Since $\mathcal{J}$ is an isometry (in fact, unitary), for each $i=1,\ldots,p$, it follows that
$\begin{split}\langle F_{0}K_{0}+f_{0},f_{i}\rangle&=\langle\mathcal{J}(k_{0}^{1},\ldots,k_{0}^{p},1),\mathcal{J}e_{i}\rangle\\ &=\langle(k_{0}^{1},\ldots,k_{0}^{p},1),e_{i}\rangle\\ &=k_{0}^{i}(0).\end{split}$
@@
proving that $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i}$. ∎
Considering the converse half of the previous statement, it is evident that the theorem concerns the classification of closed subspaces of $H^{2}$ that are invariant under Sarason-type perturbations of the backward shift $T_{z}^{*}$. Furthermore, in the setting of Corollary 2.2, the same proof of the converse part also shows that $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i}$. The following alternative can be considered a variation of Corollary 2.2:
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ such that
@@
where $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Moreover, $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i}$.
@@ 4. Perturbations of $T_{z}$
Let $T\in\mathcal{B}(\mathcal{H})$ and let $\mathcal{M}$ is a closed subspace of $\mathcal{H}$. In the introductory section, we already pointed out the elementary fact that $T\mathcal{M}\subseteq\mathcal{M}$ if and only if $T^{*}\mathcal{M}^{\perp}\subseteq\mathcal{M}^{\perp}$. However, if $\mathcal{M}$ is explicit, then it need not be the case for $\mathcal{M}^{\perp}$. For instance, the concrete representations of finite-rank perturbations of $T_{z}^{*}$-invariant subspaces do not seem to yield concrete representations of finite-rank perturbations of $T_{z}$-invariant subspaces. In this section, we examine a class of perturbations of $T_{z}$, and obtain explicit representations of invariant subspaces of the perturbations. As hinted above, the primary method will involve exploiting the adjoint of the operator used in Theorem 2.1.
@@
$\mathcal{M}^{\perp}=[\varphi_{1},\ldots,\varphi_{p},1]\mathcal{K}_{\Psi}.$
Let us assume that $g\in H^{2}$. Then $g\in\mathcal{M}^{\perp}$ if and only if $\langle g,f\rangle=0$ for all $f\in\mathcal{M}^{\perp}$. By the above representation of $\mathcal{M}^{\perp}$, we consider an arbitrary $f$ from $\mathcal{M}^{\perp}$ as $f=\sum_{i=1}^{p}\varphi_{i}k_{i}+k_{p+1}$ for some (again, arbitrary) $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$, and compute
@@
To formulate a more convenient converse statement, we will now concentrate on Sarason-type perturbations of the forward shift (see (1.6)). Recall, for each $\varphi\in L^{\infty}$, the Toeplitz operator $T_{\varphi}$ on $H^{2}$ is defined by
$T_{\varphi}f=P_{H^{2}}(\varphi f),$
for all $f\in H^{2}$. We will need the well-known algebraic characterization of Toeplitz operators: Let $T\in\mathcal{B}(H^{2})$. Then $T$ is Toeplitz if and only if
@@
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. Assume that $\{\varphi_{i}\}_{i=1}^{p}\subseteq H^{\infty}$ is an orthonormal basis for $\mathcal{L}$. If $\mathcal{M}$ is invariant under $T_{z}-\sum_{i=1}^{m}T_{z}^{*}u_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
@@
Now let us reexamine the discussion that takes place at the end of Section 3. In particular, we continue the results described in Corollary 3.2. Along this line, we note that the subspace $\mathcal{M}$ above is also invariant under $T_{z}-\sum_{i=1}^{m}\varphi_{i}\otimes T_{z}^{*}\varphi_{i}$. To confirm this, we proceed as follows: Pick $f\in\mathcal{M}^{\perp}$. By Theorem 2.1, we have
@@
where $F=\sum_{i=0}^{\infty}A_{i}z^{i}$, $A_{i}\in\mathbb{C}^{p}$ for all $i$, and $f_{0}=\sum_{i=0}^{\infty}g_{i}(0)z^{i}$. Then
$\begin{split}(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})f&=(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})(F_{0}F+f_{0})\\ &=T_{z}^{*}(F_{0}F)+T_{z}^{*}f_{0}-\sum_{i=1}^{m}\langle F_{0}F+f_{0},\varphi_{i}\rangle T_{z}^{*}\varphi_{i}.\end{split}$
@@
$(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})f=F_{0}(T_{z}^{*}F)+T_{z}^{*}f_{0}\in\mathcal{M}^{\perp},$
which proves that $\mathcal{M}^{\perp}$ is invariant under the Sarason-type perturbation $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}\varphi_{i}\otimes\varphi_{i}$.
5. Almost invariant subpsaces
@@
Recall also that the defect of $\mathcal{M}$ is the lowest possible dimension of such a space $\mathcal{F}$. This section aims to link almost invariant subspaces of $T_{z}^{*}$ with the invariant subspaces of finite-rank perturbations for $T_{z}^{*}$. The link is quite natural and could even be considered a simple observation. However, this link in particular gives a new insight into the theory of almost invariant subspaces and unifies the perturbation theory from an invariant subspace stance. We present the result in finer generality. We will assume $\mathcal{H}$ is a Hilbert space, and $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets in $\mathcal{H}$.
@@
Let $T\in\mathcal{B}(\mathcal{H})$ and let $\mathcal{M}\subseteq\mathcal{H}$ be a closed subspace. If $\mathcal{M}\subseteq\mathcal{H}$ is invariant under $T-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then $\mathcal{M}$ is almost invariant under $T$ with defect at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T$ with defect $m$, then $\mathcal{M}$ is invariant under $T-\sum_{i=1}^{m}f_{i}\otimes T^{*}f_{i}$, where $\{f_{i}\}_{i=1}^{m}$ is an orthonormal basis for the defect space.
@@
and hence, $\mathcal{M}$ is almost invariant under $T$ with defect space $\subseteq\text{span}\{v_{i}:i=1,\ldots,m\}$. This also says, in particular, that the defect dimension is at most $m$. For the converse direction, assume that $\mathcal{M}$ is almost invariant under $T$ with defect $m$ and that $\mathcal{F}$ is the $m$-dimensional defect space having an orthonormal basis, say $\{f_{i}:1\leq i\leq m\}$. Therefore, we have $T\mathcal{M}\subseteq\mathcal{M}\oplus\mathcal{F}$. Fix $f\in\mathcal{M}$. Since $Tf\in\mathcal{M}\oplus\mathcal{F}$, there exist $f_{\mathcal{M}}\in\mathcal{M}$ and $f_{\mathcal{F}}\in\mathcal{F}$ such that
@@
Also note that $f_{\mathcal{F}}=\sum_{i=1}^{m}\langle f_{\mathcal{F}},f_{i}\rangle f_{i}$. Therefore, we have
$\begin{split}(T-\sum_{i=1}^{m}f_{i}\otimes T^{*}f_{i})f&=f-\sum_{i=1}^{m}\langle f,T^{*}f_{i}\rangle f_{i}\\ &=f-\sum_{i=1}^{m}\langle f_{\mathcal{M}}\oplus f_{\mathcal{F}},f_{i}\rangle f_{i}\\ &=f-\sum_{i=1}^{m}\langle f_{\mathcal{F}},f_{i}\rangle f_{i}\\ &=(f_{\mathcal{M}}\oplus f_{\mathcal{F}})-f_{\mathcal{F}}\\ &=f_{\mathcal{M}}\\ &\in\mathcal{M}.\end{split}$
@@
In particular, we have the following in the setting of $T^{*}_{z}$: Let $\mathcal{M}$ be a closed subspace of $H^{2}$, and let $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ be orthonormal sets in $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect dimension at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect dimension $m$, then $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}f_{i}\otimes T_{z}f_{i}$, where $\{f_{i}:1\leq i\leq m\}$ is an orthonormal basis of the $m$ dimensional defect space.
In Theorem 3.1, we have already classified the invariant subspaces of finite rank perturbations of $T_{z}^{*}$ of the above type. As a result, the above observation, along with Theorem 3.1, easily implies the following classification of almost invariant subspaces of $T_{z}^{*}$:
@@
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. If $\mathcal{M}$ is almost invariant under $T_{z}^{*}$, then there exist natural numbers $p^{\prime}<p$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$, and a unitary operator $\mathcal{J}:\mathcal{K}_{\Psi}\to\mathcal{M}$ such that
@@
where $\{f_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}(u_{1}),\ldots,P_{\mathcal{M}}(u_{m})\}$, and $\{u_{1},\ldots,u_{m}\}$ is an orthonormal basis for the defect space, and
@@
Conversely, if $\mathcal{M}$ has the representation (5.1) for some orthonormal set of vectors $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ with $\mathcal{K}_{\Psi}$ as in (5.3), and a unitary $\mathcal{J}$ as in (5.2), then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$.
A similar classification of almost invariant subspaces of $T_{z}^{*}$ was earlier obtained by Chalendar, Gallardo-Gutiérrez, and Partington [4, Theorem 3.4]. However, the present result differs substantially. First, it connects with a topic that appears to have a different flavor, namely finite-rank perturbations of $T_{z}^{*}$; second, this comes as a byproduct of finite-rank perturbations of $T_{z}^{*}$. Third, the connection with the perturbation theory made the representations of almost invariant subspaces more explicit. For instance, the representation of the model space $\mathcal{K}_{\Psi}$ is a new addition compared to the one derived in [4, Theorem 3.4].

### v2

@@ 3. Perturbations of $T_{z}^{*}$
The setting of Theorem 2.1 is exceedingly broad, embracing invariant subspaces of finite-rank perturbations of $T_{\varphi}^{*}$. This one-sided result fits Hitt and Sarason’s perspectives. Indeed, their results were only one-directional. In this section, we depart from this pattern and try to explain a converse direction. Our converse, however, makes better sense for the particular case of the special inner function: $\varphi(z)=z$, $z\in\mathbb{D}$.
Thus, we will follow the setting of Corollary 2.2 and subsequently keep on assuming that $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets of vectors in $H^{2}$. To address the reverse direction, we first note a consequence of Corollary 2.2: There exists a unitary operator $\mathcal{J}:\mathcal{K}\to\mathcal{M}$, such that
@@
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$. This readily follows from the fact that an element $(F,f_{0})\in\mathcal{K}_{\Psi}$ can be written as $(F,f_{0})=(k_{1},\ldots,k_{p},k_{p+1})$ (here we are rewriting $f_{0}=k_{p+1}$ for notational simplicity), so that
@@
We also have the identity that
$\|F_{0}F+f_{0}\|^{2}=\|F\|^{2}+\|f_{0}\|^{2}.$
We fix the rank-$m$ perturbation $T_{z}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, and set
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}u_{i}\otimes u_{i}$, then there exist a natural number $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
@@
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F_{0}F+f_{0}\|^{2}=\|{F}\|^{2}+\|{f_{0}}\|^{2}\},$ (3.3)
@@
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$. Conversely, if $\mathcal{M}$ has the representation (3.2) for some orthonormal set of vectors $\{f_{i}\}_{i=1}^{p}$ in $\mathcal{M}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ with $\mathcal{K}_{\Psi}$ as in (3.3), and a unitary $\mathcal{J}$ as in (3.4), then $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}f_{i}\otimes f_{i}$.
@@
=$\mathcal{M}=\{f\in H^{2}:f=F_{0}F+f_{0}\text{ for some }(F,f_{0})\in\mathcal{K}_{\Psi}\text{ with }\|f\|^{2}=\|F\|^{2}+\|{f_{0}}\|^{2}\}.$
@@
$(T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}f_{i}\otimes f_{i})f$ $=[T_{z}^{*}-\sum_{i=1}^{p}\langle\cdot,f_{i}\rangle T_{z}^{*}f_{i}](F_{0}F+f_{0})$
@@
for all $i=1,\ldots,p$. This conclusion arises immediately from the identity (3.1). Since $\mathcal{J}$ is an isometry (in fact, unitary), for each $i=1,\ldots,p$, it follows that
$\begin{split}\langle F_{0}K_{0}+f_{0},f_{i}\rangle&=\langle\mathcal{J}(k_{0}^{1},\ldots,k_{0}^{p},f_{0}),\mathcal{J}e_{i}\rangle\\ &=\langle(k_{0}^{1},\ldots,k_{0}^{p},f_{0}),e_{i}\rangle\\ &=k_{0}^{i}(0).\end{split}$
@@
proving that $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}f_{i}\otimes f_{i}$. Take note that in the above, we have used the same notation $T_{z}^{*}$ to denote the operator $(T_{z}^{*}\otimes I_{\mathbb{C}^{p}})$ on $H^{2}_{\mathbb{C}^{p}}$. ∎
Considering the converse half of the previous statement, it is evident that the theorem concerns the classification of closed subspaces of $H^{2}$ that are invariant under Sarason-type perturbations of $T_{z}^{*}$. Furthermore, in the setting of Corollary 2.2, the same proof of the converse part also shows that $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}f_{i}\otimes f_{i}$. The following alternative can be considered a variation of Corollary 2.2:
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
@@
where $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Moreover, $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}f_{i}\otimes f_{i}$.
@@ 4. Perturbations of $T_{z}$
Let $T\in\mathcal{B}(\mathcal{H})$ and let $\mathcal{M}$ be a closed subspace of $\mathcal{H}$. In the introductory section, we already pointed out the elementary fact that
$T\mathcal{M}\subseteq\mathcal{M},$
if and only if
$T^{*}\mathcal{M}^{\perp}\subseteq\mathcal{M}^{\perp}.$
However, if $\mathcal{M}$ is explicit, then it need not be the case for $\mathcal{M}^{\perp}$. For instance, the concrete representations of finite-rank perturbations of $T_{z}^{*}$-invariant subspaces do not seem to yield concrete representations of finite-rank perturbations of $T_{z}$-invariant subspaces. In this section, we examine a class of perturbations of $T_{z}$, and obtain explicit representations of invariant subspaces of the perturbations. As hinted above, the primary method will involve exploiting the adjoint of the perturbation used in Theorem 2.1.
@@
$\mathcal{M}^{\perp}=[\varphi_{1},\ldots,\varphi_{p},1]_{1\times(p+1)}\mathcal{K}_{\Psi}.$
Let us assume that $g\in H^{2}$. Then $g\in\mathcal{M}^{\perp}$ if and only if $\langle g,f\rangle=0$ for all $f\in\mathcal{M}^{\perp}$. By the above representation of $\mathcal{M}^{\perp}$, we consider an arbitrary $f$ from $\mathcal{M}^{\perp}$ as $f=\sum_{i=1}^{p}\varphi_{i}k_{i}+k_{p+1}$ for some $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$, and compute
@@
To formulate a more convenient converse statement, we will now concentrate on Sarason-type perturbations of the forward shift (see (1.7)). Recall the definition of Toeplitz operators (see (1.6)). The well-known algebraic characterizations of Toeplitz operators state the following: Let $T\in\mathcal{B}(H^{2})$. Then $T$ is Toeplitz if and only if
@@
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. Assume that $\{\varphi_{i}\}_{i=1}^{p}\subseteq H^{\infty}$ is an orthonormal basis for $\mathcal{L}$. If $\mathcal{M}$ is invariant under $T_{z}-\sum_{i=1}^{m}T_{z}^{*}v_{i}\otimes v_{i}$, then there exist $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
@@
Now let us reexamine the discussion that took place at the end of Section 3. In particular, we continue the results described in Corollary 3.2. Along this line, we note that the subspace $\mathcal{M}$ above is also invariant under $T_{z}-\sum_{i=1}^{p}\varphi_{i}\otimes T_{z}^{*}\varphi_{i}$. To confirm this, we proceed as follows: Pick $f\in\mathcal{M}^{\perp}$. By Corollary 3.2, we have
@@
where $F=\sum_{i=0}^{\infty}A_{i}z^{i}$, $A_{i}\in\mathbb{C}^{p}$ for all $i$, and $f_{0}=\sum_{i=0}^{\infty}g_{i+1}(0)z^{i}$. Then
$\begin{split}(T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})f&=(T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})(F_{0}F+f_{0})\\ &=T_{z}^{*}(F_{0}F)+T_{z}^{*}f_{0}-\sum_{i=1}^{p}\langle F_{0}F+f_{0},\varphi_{i}\rangle T_{z}^{*}\varphi_{i}.\end{split}$
@@
$(T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})f=F_{0}(T_{z}^{*}F)+T_{z}^{*}f_{0}\in\mathcal{M}^{\perp},$
proves that $\mathcal{M}^{\perp}$ is invariant under the Sarason-type perturbation $T_{z}^{*}-\sum_{i=1}^{p}T_{z}^{*}\varphi_{i}\otimes\varphi_{i}$.
5. Almost invariant subspaces
@@
Recall also that the defect of $\mathcal{M}$ is the lowest possible dimension of such a space $\mathcal{F}$. This section aims to link almost invariant subspaces of $T_{z}^{*}$ with the invariant subspaces of finite-rank perturbations for $T_{z}^{*}$. The link is quite natural and could even be considered a simple observation. However, this in particular gives a new insight into the theory of almost invariant subspaces and unifies the perturbation theory from an invariant subspace stance. We present the result in finer generality. We will assume $\mathcal{H}$ is a Hilbert space, and, as usual, $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets in $\mathcal{H}$.
@@
Let $T\in\mathcal{B}(\mathcal{H})$ and let $\mathcal{M}\subseteq\mathcal{H}$ be a closed subspace. If $\mathcal{M}$ is invariant under $T-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then $\mathcal{M}$ is almost invariant under $T$ with defect at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T$ with defect $m$, then $\mathcal{M}$ is invariant under $T-\sum_{i=1}^{m}f_{i}\otimes T^{*}f_{i}$, where $\{f_{i}\}_{i=1}^{m}$ is an orthonormal basis for the defect space.
@@
and hence, $\mathcal{M}$ is almost invariant under $T$ with defect space $\mathcal{F}$ such that
$\mathcal{F}\subseteq\text{span}\{v_{i}:i=1,\ldots,m\}.$
This also says, in particular, that the $\dim\mathcal{F}$ is at most $m$. For the converse direction, assume that $\mathcal{M}$ is almost invariant under $T$ with defect $m$ and that $\mathcal{F}$ is the $m$-dimensional defect space having an orthonormal basis, say $\{f_{i}:1\leq i\leq m\}$. Therefore, we have
$T\mathcal{M}\subseteq\mathcal{M}\oplus\mathcal{F}.$
Fix $f\in\mathcal{M}$. Since $Tf\in\mathcal{M}\oplus\mathcal{F}$, there exist $f_{\mathcal{M}}\in\mathcal{M}$ and $f_{\mathcal{F}}\in\mathcal{F}$ such that
@@
Therefore, we have
$\begin{split}(T-\sum_{i=1}^{m}f_{i}\otimes T^{*}f_{i})f&=Tf-\sum_{i=1}^{m}\langle f,T^{*}f_{i}\rangle f_{i}\\ &=Tf-\sum_{i=1}^{m}\langle f_{\mathcal{M}}\oplus f_{\mathcal{F}},f_{i}\rangle f_{i}\\ &=Tf-\sum_{i=1}^{m}\langle f_{\mathcal{F}},f_{i}\rangle f_{i}.\end{split}$
Also note that
$f_{\mathcal{F}}=\sum_{i=1}^{m}\langle f_{\mathcal{F}},f_{i}\rangle f_{i},$
and consequently
$\begin{split}(T-\sum_{i=1}^{m}f_{i}\otimes T^{*}f_{i})f&=Tf-\sum_{i=1}^{m}\langle f_{\mathcal{F}},f_{i}\rangle f_{i}\\ &=(f_{\mathcal{M}}\oplus f_{\mathcal{F}})-f_{\mathcal{F}}\\ &=f_{\mathcal{M}}\\ &\in\mathcal{M}.\end{split}$
@@
In particular, we have the following in the setting of $T^{*}_{z}$: Let $\mathcal{M}$ be a closed subspace of $H^{2}$, and let $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ be orthonormal sets in $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect $m$, then $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}f_{i}\otimes T_{z}f_{i}$, where $\{f_{i}:1\leq i\leq m\}$ is an orthonormal basis of the defect space.
In Theorem 3.1, we have already classified the invariant subspaces of finite-rank perturbations of $T_{z}^{*}$ of the above type. As a result, the above observation, along with Theorem 3.1, easily implies the classification of almost invariant subspaces of $T_{z}^{*}$:
@@
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. If $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect $m$, then there exist a natural number $p^{\prime}\leq p+1$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$, and a unitary operator $\mathcal{J}:\mathcal{K}_{\Psi}\to\mathcal{M}$ such that
@@
where $F_{0}=[f_{1},\ldots,f_{p}]$, $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}(T_{z}u_{i})\}_{i=1}^{m}$, and $\{u_{i}\}_{i=1}^{m}$ is an orthonormal basis for the defect space, and
@@
Conversely, if $\mathcal{M}$ has the representation (5.1) for some orthonormal set of vectors $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{M}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ with $\mathcal{K}_{\Psi}$ as in (5.3), and a unitary $\mathcal{J}$ as in (5.2), then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$.
We remark that Chalendar, Gallardo-Gutiérrez, and Partington previously classified almost invariant subspaces of $T_{z}^{*}$ [4, Theorem 3.4]. However, the present result differs substantially. First, it connects with a topic that appears to have a different flavor, namely finite-rank perturbations of $T_{z}^{*}$; second, this comes as a byproduct of finite-rank perturbations of $T_{z}^{*}$. Third, the connection with the perturbation theory made the representations of almost invariant subspaces more explicit. For instance, the representation of the model space $\mathcal{K}_{\Psi}$ is a new addition compared to the one derived in [4, Theorem 3.4].